\begin{thebibliography}{}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Anderson et~al.(2016)Anderson, Fernando, Johnson, and Gould}]{spice}
Peter Anderson, Basura Fernando, Mark Johnson, and Stephen Gould. 2016.
\newblock Spice: Semantic propositional image caption evaluation.
\newblock In {\em Computer Vision--ECCV 2016: 14th European Conference,
  Amsterdam, The Netherlands, October 11-14, 2016, Proceedings, Part V 14\/}.
  Springer, pages 382--398.

\bibitem[{Banerjee and Lavie(2005)}]{meteor}
Satanjeev Banerjee and Alon Lavie. 2005.
\newblock Meteor: An automatic metric for mt evaluation with improved
  correlation with human judgments.
\newblock In {\em Proceedings of the acl workshop on intrinsic and extrinsic
  evaluation measures for machine translation and/or summarization\/}. pages
  65--72.

\bibitem[{Bao et~al.(2020)Bao, Dong, Wei, Wang, Yang, Liu, Wang, Gao, Piao,
  Zhou et~al.}]{unilmv2:2020}
Hangbo Bao, Li~Dong, Furu Wei, Wenhui Wang, Nan Yang, Xiaodong Liu, Yu~Wang,
  Jianfeng Gao, Songhao Piao, Ming Zhou, et~al. 2020.
\newblock Unilmv2: Pseudo-masked language models for unified language model
  pre-training.
\newblock In {\em International conference on machine learning\/}. PMLR, pages
  642--652.

\bibitem[{Barzilay and Lapata(2005)}]{barzilay-lapata-2005-modeling}
Regina Barzilay and Mirella Lapata. 2005.
\newblock \href{https://doi.org/10.3115/1219840.1219858}{Modeling local
  coherence: An entity-based approach}.
\newblock In {\em Proceedings of the 43rd Annual Meeting of the Association for
  Computational Linguistics ({ACL}{'}05)\/}. Association for Computational
  Linguistics, Ann Arbor, Michigan, pages 141--148.
\newblock
  \href{https://doi.org/10.3115/1219840.1219858}{https://doi.org/10.3115/1219840.1219858}.

\bibitem[{Blodgett et~al.(2021)Blodgett, Lopez, Olteanu, Sim, and
  Wallach}]{blodgett2021stereotyping}
Su~Lin Blodgett, Gilsinia Lopez, Alexandra Olteanu, Robert Sim, and Hanna
  Wallach. 2021.
\newblock Stereotyping norwegian salmon: An inventory of pitfalls in fairness
  benchmark datasets.
\newblock In {\em Proceedings of the 59th Annual Meeting of the Association for
  Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)\/}. pages 1004--1015.

\bibitem[{Bollegala et~al.(2005)Bollegala, Okazaki, and
  Ishizuka}]{bollegala-etal-2005-machine}
Danushka Bollegala, Naoaki Okazaki, and Mitsuru Ishizuka. 2005.
\newblock \href{https://doi.org/10.1007/11562214\_55}{A machine learning
  approach to sentence ordering for multidocument summarization and its
  evaluation}.
\newblock In {\em Second International Joint Conference on Natural Language
  Processing: Full Papers\/}.
\newblock
  \href{https://doi.org/10.1007/11562214\_55}{https://doi.org/10.1007/11562214\_55}.

\bibitem[{Bollegala et~al.(2006)Bollegala, Okazaki, and
  Ishizuka}]{bollegala-etal-2006-bottom}
Danushka Bollegala, Naoaki Okazaki, and Mitsuru Ishizuka. 2006.
\newblock \href{https://doi.org/10.3115/1220175.1220224}{A bottom-up approach
  to sentence ordering for multi-document summarization}.
\newblock In {\em Proceedings of the 21st International Conference on
  Computational Linguistics and 44th Annual Meeting of the Association for
  Computational Linguistics\/}. Association for Computational Linguistics,
  Sydney, Australia, pages 385--392.
\newblock
  \href{https://doi.org/10.3115/1220175.1220224}{https://doi.org/10.3115/1220175.1220224}.

\bibitem[{Bollegala et~al.(2012)Bollegala, Okazaki, and
  Ishizuka}]{Bollegala:IS:2012}
Danushka Bollegala, Naoaki Okazaki, and Mitsuru Ishizuka. 2012.
\newblock A preference learning approach to sentence ordering for
  multi-document summarization.
\newblock {\em Information Sciences\/} 217:78 -- 95.

\bibitem[{Brown et~al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal,
  Neelakantan, Shyam, Sastry, Askell et~al.}]{GPT3:2020}
Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared~D Kaplan, Prafulla
  Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell,
  et~al. 2020.
\newblock Language models are few-shot learners.
\newblock {\em Advances in neural information processing systems\/}
  33:1877--1901.

\bibitem[{Chen et~al.(2023)Chen, Xu, Yan, Zhang, Huang, Si, and
  Zhang}]{chen:2023}
Qianglong Chen, Guohai Xu, Ming Yan, Ji~Zhang, Fei Huang, Luo Si, and Yin
  Zhang. 2023.
\newblock \href{https://doi.org/10.18653/v1/2023.findings-acl.835}{Distinguish
  before answer: Generating contrastive explanation as knowledge for
  commonsense question answering}.
\newblock In {\em Findings of the Association for Computational Linguistics:
  ACL 2023\/}. Association for Computational Linguistics, Toronto, Canada,
  pages 13207--13224.
\newblock
  \href{https://doi.org/10.18653/v1/2023.findings-acl.835}{https://doi.org/10.18653/v1/2023.findings-acl.835}.

\bibitem[{Fan et~al.(2020)Fan, Gong, Wei, Wang, Huang, Jiao, Huang, Duan, and
  Zhang}]{Ekibart:2020}
Zhihao Fan, Yeyun Gong, Zhongyu Wei, Siyuan Wang, Yameng Huang, Jian Jiao,
  Xuan-Jing Huang, Nan Duan, and Ruofei Zhang. 2020.
\newblock An enhanced knowledge injection model for commonsense generation.
\newblock In {\em Proceedings of the 28th International Conference on
  Computational Linguistics\/}. pages 2014--2025.

\bibitem[{Gardent et~al.(2017)Gardent, Shimorina, Narayan, and
  Perez-Beltrachini}]{Gardent:2017}
Claire Gardent, Anastasia Shimorina, Shashi Narayan, and Laura
  Perez-Beltrachini. 2017.
\newblock The {W}eb{NLG} challenge: Generating text from {RDF} data.
\newblock In {\em Proceedings of the 10th International Conference on Natural
  Language Generation\/}. Association for Computational Linguistics, Santiago
  de Compostela, Spain, pages 124--133.

\bibitem[{Gatt and Krahmer(2018)}]{Gatt:2018}
Albert Gatt and Emiel Krahmer. 2018.
\newblock Survey of the state of the art in natural language generation: Core
  tasks, applications and evaluation.
\newblock {\em J. Artif. Int. Res.\/} 61(1):65--170.

\bibitem[{Guti{\'e}rrez et~al.(2022)Guti{\'e}rrez, McNeal, Washington, Chen,
  Li, Sun, and Su}]{gutierrez:2022}
Bernal~Jim{\'e}nez Guti{\'e}rrez, Nikolas McNeal, Clay Washington, You Chen,
  Lang Li, Huan Sun, and Yu~Su. 2022.
\newblock Thinking about gpt-3 in-context learning for biomedical ie? think
  again.
\newblock {\em arXiv preprint arXiv:2203.08410\/} .

\bibitem[{Hoyle et~al.(2021)Hoyle, Marasovi{\'c}, and Smith}]{hoyle:2021}
Alexander~Miserlis Hoyle, Ana Marasovi{\'c}, and Noah~A Smith. 2021.
\newblock Promoting graph awareness in linearized graph-to-text generation.
\newblock In {\em Findings of the Association for Computational Linguistics:
  ACL-IJCNLP 2021\/}. pages 944--956.

\bibitem[{Huang et~al.(2023)Huang, Gao, Li, Yang, Song, Zhang, Zhu, Jiang,
  Chang, and Yin}]{huang:2023}
Jie Huang, Yifan Gao, Zheng Li, Jingfeng Yang, Yangqiu Song, Chao Zhang, Zining
  Zhu, Haoming Jiang, Kevin Chen-Chuan Chang, and Bing Yin. 2023.
\newblock Ccgen: Explainable complementary concept generation in e-commerce.
\newblock {\em arXiv preprint arXiv:2305.11480\/} .

\bibitem[{Kingma and Ba(2015)}]{Adam}
Diederik~P. Kingma and Jimmy Ba. 2015.
\newblock Adam: {A} method for stochastic optimization.
\newblock In Yoshua Bengio and Yann LeCun, editors, {\em 3rd International
  Conference on Learning Representations, {ICLR} 2015, San Diego, CA, USA, May
  7-9, 2015, Conference Track Proceedings\/}.

\bibitem[{Lewis et~al.(2019)}]{BART:2019}
Michael Lewis et~al. 2019.
\newblock Bart: Denoising sequence-to-sequence pre-training for natural
  language generation, translation, and comprehension.
\newblock {\em meeting of the association for computational linguistics\/} .

\bibitem[{Lin et~al.(2020)Lin, Zhou, Shen, Zhou, Bhagavatula, Choi, and
  Ren}]{CommonGen}
Bill~Yuchen Lin, Wangchunshu Zhou, Ming Shen, Pei Zhou, Chandra Bhagavatula,
  Yejin Choi, and Xiang Ren. 2020.
\newblock {C}ommon{G}en: A constrained text generation challenge for generative
  commonsense reasoning.
\newblock In {\em Findings of the Association for Computational Linguistics:
  EMNLP 2020\/}. Association for Computational Linguistics, Online, pages
  1823--1840.

\bibitem[{Lin(2004)}]{rouge}
Chin-Yew Lin. 2004.
\newblock Rouge: A package for automatic evaluation of summaries.
\newblock In {\em Text summarization branches out\/}. pages 74--81.

\bibitem[{Liu et~al.(2022)Liu, Liu, Yang, Zhang, Ding, Yao, Luo, Zhang, and
  Su}]{Kgr4:2022}
Xin Liu, Dayiheng Liu, Baosong Yang, Haibo Zhang, Junwei Ding, Wenqing Yao,
  Weihua Luo, Haiying Zhang, and Jinsong Su. 2022.
\newblock Kgr4: Retrieval, retrospect, refine and rethink for commonsense
  generation.
\newblock In {\em Proceedings of the AAAI Conference on Artificial
  Intelligence\/}. volume~36, pages 11029--11037.

\bibitem[{Liu et~al.(2021)Liu, Wan, He, Peng, and Yu}]{kgbart:2021}
Ye~Liu, Yao Wan, Lifang He, Hao Peng, and Philip~S Yu. 2021.
\newblock Kg-bart: Knowledge graph-augmented bart for generative commonsense
  reasoning.
\newblock In {\em Proceedings of the AAAI Conference on Artificial
  Intelligence\/}. volume~35, pages 6418--6425.

\bibitem[{Lu et~al.(2021)Lu, West, Zellers, Le~Bras, Bhagavatula, and
  Choi}]{Neurologic:2021}
Ximing Lu, Peter West, Rowan Zellers, Ronan Le~Bras, Chandra Bhagavatula, and
  Yejin Choi. 2021.
\newblock Neurologic decoding:(un) supervised neural text generation with
  predicate logic constraints.
\newblock In {\em Proceedings of the 2021 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies\/}. pages 4288--4299.

\bibitem[{Moryossef et~al.(2019)Moryossef, Goldberg, and
  Dagan}]{Moryossef:2019}
Amit Moryossef, Yoav Goldberg, and Ido Dagan. 2019.
\newblock {S}tep-by-step: {S}eparating planning from realization in neural
  data-to-text generation.
\newblock In {\em Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)\/}. Association for
  Computational Linguistics, Minneapolis, Minnesota, pages 2267--2277.

\bibitem[{Mostafazadeh et~al.(2016)Mostafazadeh, Chambers, He, Parikh, Batra,
  Vanderwende, Kohli, and Allen}]{rocstories}
Nasrin Mostafazadeh, Nathanael Chambers, Xiaodong He, Devi Parikh, Dhruv Batra,
  Lucy Vanderwende, Pushmeet Kohli, and James Allen. 2016.
\newblock A corpus and cloze evaluation for deeper understanding of commonsense
  stories.
\newblock In {\em Proceedings of the 2016 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies\/}. pages 839--849.

\bibitem[{Ou et~al.(2022)Ou, Zhang, and Zhang}]{Ou:2022}
Zebin Ou, Meishan Zhang, and Yue Zhang. 2022.
\newblock On the role of pre-trained language models in word ordering: A case
  study with bart.
\newblock In {\em Proceedings of the 29th International Conference on
  Computational Linguistics\/}. pages 6516--6529.

\bibitem[{Papineni et~al.(2002)Papineni, Roukos, Ward, and Zhu}]{bleu}
Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002.
\newblock Bleu: a method for automatic evaluation of machine translation.
\newblock In {\em Proceedings of the 40th annual meeting of the Association for
  Computational Linguistics\/}. pages 311--318.

\bibitem[{Pennington et~al.(2014)Pennington, Socher, and Manning}]{glove}
Jeffrey Pennington, Richard Socher, and Christopher~D Manning. 2014.
\newblock Glove: Global vectors for word representation.
\newblock In {\em Proceedings of the 2014 conference on empirical methods in
  natural language processing (EMNLP)\/}. pages 1532--1543.

\bibitem[{Puzikov and Gurevych(2018)}]{Puzikov:2018}
Yevgeniy Puzikov and Iryna Gurevych. 2018.
\newblock {E}2{E} {NLG} challenge: Neural models vs. templates.
\newblock In {\em Proceedings of the 11th International Conference on Natural
  Language Generation\/}. Association for Computational Linguistics, Tilburg
  University, The Netherlands, pages 463--471.

\bibitem[{Raffel et~al.(2019)Raffel, Shazeer, Roberts, Lee, Narang, Matena,
  Zhou, Li, and Liu}]{T5:2019}
Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael
  Matena, Yanqi Zhou, Wei Li, and Peter~J. Liu. 2019.
\newblock Exploring the limits of transfer learning with a unified text-to-text
  transformer.
\newblock {\em Journal of Machine Learning Research\/} .

\bibitem[{Reiter and Dale(2000)}]{Reiter:2000}
Ehud Reiter and Robert Dale. 2000.
\newblock {\em Building Natural Language Generation Systems\/}.
\newblock Cambridge University Press.

\bibitem[{Sharma et~al.(2021)Sharma, Dey, and Sinha}]{sharma:2021}
Shanya Sharma, Manan Dey, and Koustuv Sinha. 2021.
\newblock \href{https://openreview.net/forum?id=bnuU0PzXl0-}{Evaluating gender
  bias in natural language inference}.
\newblock
  \href{https://openreview.net/forum?id=bnuU0PzXl0-}{https://openreview.net/forum?id=bnuU0PzXl0-}.

\bibitem[{Sutskever et~al.(2014)Sutskever, Vinyals, and
  Le}]{Sutskever:NIPS:2014}
Ilya Sutskever, Oriol Vinyals, and Quoc~V. Le. 2014.
\newblock Sequence to sequence learning with neural networks.
\newblock In {\em NIPS'14\/}.

\bibitem[{Thoppilan et~al.(2022)Thoppilan, De~Freitas, Hall, Shazeer,
  Kulshreshtha, Cheng, Jin, Bos, Baker, Du et~al.}]{lamda:2022}
Romal Thoppilan, Daniel De~Freitas, Jamie Hall, Noam Shazeer, Apoorv
  Kulshreshtha, Heng-Tze Cheng, Alicia Jin, Taylor Bos, Leslie Baker, Yu~Du,
  et~al. 2022.
\newblock Lamda: Language models for dialog applications.
\newblock {\em arXiv preprint arXiv:2201.08239\/} .

\bibitem[{Vedantam et~al.(2015)Vedantam, Lawrence~Zitnick, and Parikh}]{cider}
Ramakrishna Vedantam, C~Lawrence~Zitnick, and Devi Parikh. 2015.
\newblock Cider: Consensus-based image description evaluation.
\newblock In {\em Proceedings of the IEEE conference on computer vision and
  pattern recognition\/}. pages 4566--4575.

\bibitem[{Wei et~al.(2022)Wei, Wang, Schuurmans, Bosma, Xia, Chi, Le, Zhou
  et~al.}]{wei:2022}
Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed~H Chi,
  Quoc~V Le, Denny Zhou, et~al. 2022.
\newblock Chain-of-thought prompting elicits reasoning in large language
  models.
\newblock In {\em Advances in Neural Information Processing Systems\/}.

\bibitem[{Wolf et~al.(2020)Wolf, Debut, Sanh, Chaumond, Delangue, Moi, Cistac,
  Rault, Louf, Funtowicz, Davison, Shleifer, von Platen, Ma, Jernite, Plu, Xu,
  Scao, Gugger, Drame, Lhoest, and Rush}]{transformers:2020}
Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue,
  Anthony Moi, Pierric Cistac, Tim Rault, Rémi Louf, Morgan Funtowicz, Joe
  Davison, Sam Shleifer, Patrick von Platen, Clara Ma, Yacine Jernite, Julien
  Plu, Canwen Xu, Teven~Le Scao, Sylvain Gugger, Mariama Drame, Quentin Lhoest,
  and Alexander~M. Rush. 2020.
\newblock
  \href{https://www.aclweb.org/anthology/2020.emnlp-demos.6}{Transformers:
  State-of-the-art natural language processing}.
\newblock In {\em Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing: System Demonstrations\/}. Association for
  Computational Linguistics, Online, pages 38--45.
\newblock
  \href{https://www.aclweb.org/anthology/2020.emnlp-demos.6}{https://www.aclweb.org/anthology/2020.emnlp-demos.6}.

\bibitem[{Yang et~al.(2023)Yang, Wang, Li, Bi, Lam, and Xu}]{Yang:2023}
Haoran Yang, Yan Wang, Piji Li, Wei Bi, Wai Lam, and Chen Xu. 2023.
\newblock Bridging the gap between pre-training and fine-tuning for commonsense
  generation.
\newblock In {\em Findings of the Association for Computational Linguistics:
  EACL 2023\/}. pages 376--383.

\bibitem[{Zhang et~al.(2023)Zhang, Zhang, Li, Zhao, Karypis, and
  Smola}]{zhang:2023}
Zhuosheng Zhang, Aston Zhang, Mu~Li, Hai Zhao, George Karypis, and Alex Smola.
  2023.
\newblock Multimodal chain-of-thought reasoning in language models.
\newblock {\em arXiv preprint arXiv:2302.00923\/} .

\bibitem[{Zhao et~al.(2022)Zhao, Brahman, Huang, and Chaturvedi}]{Zhao:2022}
Chao Zhao, Faeze Brahman, Tenghao Huang, and Snigdha Chaturvedi. 2022.
\newblock \href{https://doi.org/10.18653/v1/2022.findings-naacl.129}{Revisiting
  generative commonsense reasoning: A pre-ordering approach}.
\newblock In {\em Findings of the Association for Computational Linguistics:
  NAACL 2022\/}. Association for Computational Linguistics, Seattle, United
  States, pages 1709--1718.
\newblock
  \href{https://doi.org/10.18653/v1/2022.findings-naacl.129}{https://doi.org/10.18653/v1/2022.findings-naacl.129}.

\bibitem[{Zhao et~al.(2020)Zhao, Walker, and Chaturvedi}]{Zhao:2020}
Chao Zhao, Marilyn Walker, and Snigdha Chaturvedi. 2020.
\newblock Bridging the structural gap between encoding and decoding for
  data-to-text generation.
\newblock In {\em Proceedings of the 58th Annual Meeting of the Association for
  Computational Linguistics\/}. Association for Computational Linguistics,
  Online, pages 2481--2491.

\bibitem[{Zhou et~al.(2021)Zhou, Lee, Selvam, Lee, and Ren}]{CALM:2021}
Wangchunshu Zhou, Dong-Ho Lee, Ravi~Kiran Selvam, Seyeon Lee, and Xiang Ren.
  2021.
\newblock \href{https://openreview.net/forum?id=3k20LAiHYL2}{Pre-training
  text-to-text transformers for concept-centric common sense}.
\newblock In {\em International Conference on Learning Representations\/}.
\newblock
  \href{https://openreview.net/forum?id=3k20LAiHYL2}{https://openreview.net/forum?id=3k20LAiHYL2}.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt,a4paper]{article}
\usepackage[hyperref]{acl2023}
\usepackage{times}
\usepackage{inconsolata}
\usepackage{latexsym}
\usepackage{danudefs}
\usepackage{algorithmic}
\usepackage{algorithm}
\usepackage{color}
\usepackage{booktabs}
\usepackage{xspace}
\usepackage{url}
\usepackage{braket}
\usepackage{multirow}
\usepackage{graphicx}
\usepackage{microtype}

\usepackage[english]{babel}
\usepackage{hyperref}
\addto\captionsenglish{%
}
\addto\extrasenglish{%
}

\usepackage{acronym}
\acrodef{MLM}[MLM]{Masked Language Model}
\acrodef{SoTA}[SoTA]{state-of-the-art}

\usepackage{todonotes}

\makeatletter
\if@todonotes@disabled

\else

\fi
\makeatother

\title{Learning to Predict Concept Ordering for Common Sense Generation}

\author{Tianhui Zhang \And Danushka Bollegala \\
  University of Liverpool, Amazon \\ 
  {\tt \{tianhui.zhang, danushka, bei.peng\}@liverpool.ac.uk} \\ \And Bei Peng}

\date{}

\begin{document}
\maketitle
\begin{abstract}
Prior work has shown that the ordering in which concepts are shown to a commonsense generator plays an important role, affecting the quality of the generated sentence. 
However, it remains a challenge to determine the optimal ordering of a given set of concepts such that a natural sentence covering all the concepts could be generated from a pretrained generator.
To understand the relationship between the ordering of the input concepts and the quality of the generated sentences, we conduct a systematic study considering multiple language models (LMs) and concept ordering strategies.
We find that \texttt{BART-large} model consistently outperforms all other LMs considered in this study when fine-tuned using the ordering of concepts as they appear in CommonGen training data  as measured using multiple evaluation metrics.
Moreover, the larger GPT3-based large language models (LLMs) variants do not necessarily outperform much smaller LMs on this task, even when fine-tuned on task-specific training data.
Interestingly, human annotators significantly reorder input concept sets when manually writing sentences covering those concepts, and this ordering provides the best sentence generations independently of the LM used for the generation, outperforming a probabilistic concept ordering baseline.\footnote{Code: \url{https://github.com/TianhuiZhang/concept_ordering}}
\end{abstract}

\section{Introduction}
\label{sec:intro}


In Generative Commonsense Reasoning (GCR)~\cite{CommonGen}, the goal is to generate a natural commonsense adhering sentence that covers all of the input concepts.
GCR is a challenging natural language generation (NLG) task because it requires (a) relational reasoning using background commonsense knowledge associated with the input concepts, and (b) compositional generalisation ability to work on unseen concept combinations.
CommonGen dataset~\cite{CommonGen} was specifically developed to evaluate the ability of a generative model to produce sentences covering a given set of concepts.
For example, given the set of concepts \{\emph{ball}, \emph{batter}, \emph{pitcher}, \emph{throw}\}, a human annotator would come up with a sentence such as \emph{The \textbf{pitcher} \textbf{throw}s the \textbf{ball}, and the \textbf{batter} hits a home run!}
The input concepts (shown in boldface fonts) are ordered in a specific manner in the produced sentence such as to create a commonsense bearing sentence.

Although neural generation systems produce fluent texts when compared to template-based methods, they fall short in fluency and faithfulness to the input and do not allow control over the output structure~\cite{Puzikov:2018}.
Similar observations are made by prior work on GCR, where the input ordering of concepts has been reported to influence the quality of the generated sentences~\cite{Zhao:2022,Yang:2023}.
However, the relationship between (a) the ordering of the concepts given as the input to a neural text generation model and (b) the architecture of the underlying LM used in the generator remains unexplored, which is the focus of our study in this paper.
Although we consider the concept ordering problem in the context of GCR, we note that it has broader relevance to other tasks in NLP such as multi-document summarisation~\cite{bollegala-etal-2005-machine,bollegala-etal-2006-bottom,Bollegala:IS:2012} and textual coherence modelling~\cite{barzilay-lapata-2005-modeling}

We evaluate the quality of the sentences generated by five commonsense generation LMs: \href{https://github.com/microsoft/unilm}{\texttt{BERT-gen}}~\cite{unilmv2:2020}, \href{https://huggingface.co/facebook/bart-base}{\texttt{BART-base}}~\cite{BART:2019}, \href{https://huggingface.co/t5-base}{\texttt{T5-base}}~\citep{T5:2019}, \href{https://huggingface.co/facebook/bart-large}{\texttt{BART-large}}~\citep{BART:2019} and \href{https://huggingface.co/t5-large}{\texttt{T5-large}}~\citep{T5:2019},
where the input concepts are ordered following three different strategies as described in \autoref{sec:order-methods}.
We use seven evaluation metrics to compare the generated sentences against human-written sentences in CommonGen.
Specifically, we fine-tune each LM on the train sentences in CommonGen dataset using next token prediction as the training objective.

We find \texttt{BART-large} to consistently outperform the other LMs across several evaluation metrics when fine-tuned using the ordering of concepts as they appear in CommonGen training data.
To further compare the concept ordering in the model-generated sentence against the human-generated sentence for the same set of concepts, we use the Kendall rank correlation coefficient (i.e., Kendall's $\tau$) as an evaluation metric.
We find that there exists only a weak correlation ($\tau = 0.328$) between the ordering of concepts shown to the human annotators and the ordering in which those concepts appear in the sentences written by those annotators.
This indicates that human annotators significantly reorder the input concepts when writing sentences that convey commonsense relationships among given concepts.
Interestingly, all the generator models we compare in this paper are able to produce better quality sentences when they are presented with the input concepts in the same ordering as ordered in the sentences.

\section{Related Work}
\label{sec:relate}

\paragraph{Concept Ordering.} Concept Ordering aims to reorder the given concepts in a sequence according to their importance and inner relevance.  Recent work by \citet{Ou:2022} highlighted the effectiveness of pre-trained language models, such as BART~\cite{BART:2019}, in tasks related to concept ordering. \citet{huang:2023} fine-tune the LM to find the most complementary concepts to the given one. \citet{hoyle:2021} and \citep{Zhao:2022} show that suitable concept ordering could increase the quality of the generated sentences. In this work, we aim to explore the relation between the concept ordering and different LMs, assessing their performance in the GCR task.

\paragraph{Generative Commonsense Reasoning.} Recently, a series of works have been proposed to evaluate the commonsense reasoning quality of the model's generation. One strand of research leverages these generations as the external commonsense explanation ~\cite{chen:2023} or chain-of-thought~\cite{zhang:2023}, aiding models in various tasks, including question answering.  Another approach compares the generation with the references, examining if the model could write as natural as human, such as CommonGen~\cite{CommonGen} and ROCStories~\cite{rocstories}. \citet{kgbart:2021} and \citet{Kgr4:2022} incorporate external knowledge into pre-trained LMs to enrich the generation information. 

\paragraph{Text Generation.} Prior work on text generation from structured data such as RDF has shown that the ordering in which a set of entities shown to neural text generation models significantly influences the quality of the generated text~\cite{Moryossef:2019,Zhao:2020}.
Unlike template-based generation methods~\cite{Reiter:2000,Gatt:2018}, neural text generation methods such as Seq2Seq~\cite{Sutskever:NIPS:2014} models perform \emph{text planning} (how to order the inputs) as well as \emph{plan realisation} (how to verbalise the plan) as a single end-to-end task~\cite{Gardent:2017}.

\section{Methods}
\label{sec:task}


\subsection{Task definition}
\label{sec:task-def}
Given a set $\cX = \{x_1, x_2, \ldots, x_m\}$ of lemmatised tokens representing concepts $x_i$, the goal of the GCR task is to generate a natural and grammatically correct sentence $\vec{y} = y_1, y_2, \ldots, y_n$, with tokens (or sub-tokens) $y_j$.
Note that $\cX$ is an \emph{unordered} set of concepts, while the tokens in $\vec{y}$ are ordered.
We must use \emph{all} concepts in $\cX$ when generating $\vec{y}$.
However, we are allowed to use different morphological forms (inflections) of the concepts for this purpose.
Typically in CommonGen dataset ca. $m = 5$, whereas $n$ is determined by the generator model used to produce $\vec{y}$.
Given a set of concepts $\cX$, there exists $m!$ number of possible ordering of concepts.

Let us denote one such ordering by $\vec{x} = x_1, x_2, \ldots, x_m$.
Given $\vec{x}$, a generator parameterised by $\vec{\theta}$ produces the output sequence $\vec{y} = y_1, y_2, \ldots y_n$ according to the generation probabilities given by \eqref{eq:gen-prob}.
\begin{align}
    \label{eq:gen-prob}
    p(\vec{y} | \vec{x}; \vec{\theta}) = \prod_{i=2}^{n} p(y_i | y_{1:i-1}, \vec{x}; \vec{\theta})
\end{align}
We generate sentences $\vec{y}$ using different pre-trained generators for the same set of concepts $\cX$, ordered using different strategies as described in \autoref{sec:order-methods}.

\subsection{Concept Ordering Strategies}
\label{sec:order-methods}
We name the ordering of input concepts in a train/test instance in CommonGen as the \textbf{Original} ordering.
We propose three different strategies to re-order the Original ordering for the purpose of fine-tuning LMs using the CommonGen training instances.
In \textbf{Random} ordering, the concepts in the input set are randomly ordered.
The \textbf{Example} ordering considers the ordering of concepts as they appear in a human-written example sentence in the CommonGen train dataset.\footnote{As detailed in \autoref{sec:format}, in our preliminary experiments, we evaluated three different input formats for a concept ordering:
(a) space delimited, (b) comma delimited, and (c) space delimited with a special end-of-ordering token (i.e., [ORDERING]). 
We found the input format method (c) to perform slightly better than the other two, albeit no significant improvements were observed.}
However, this concept ordering information is not available at test time, we thus use Random Ordering of input concept sets at test time even for the LMs fine-tuned using the Example Ordering.

We induce a \textbf{Probabilistic} ordering among the concepts in a given set using transition probabilities $p(x_j|x_i)$ for ordering the concept $x_j$ after $x_i$ because we want to find a method that could be used both at training and test times and do not extract the concept ordering from pre-trained models' outputs. Therefore, inspired by Glove word embeddings~\cite{glove}, which use the co-occurrence information between words, we use the co-occurrence frequency of each concept pair inside the paths of the ConceptNet to determine the ordering in a given concept set. 
Specifically, we first perform random walks over the ConceptNet graph starting from vertices that appear in the CommonGen train sentences, limiting to a maximum path length of five concepts.
Next, we count the number of paths, $\#(x_i \rightarrow x_j)$, where $x_i$ appears before $x_j$.
The transition probabilities are estimated from the path counts as in \eqref{eq:prob}.
\begin{align}
    \label{eq:prob}
    p(x_j|x_i) = \frac{\#(x_i \rightarrow x_j)}{\#(x_i \rightarrow x_j) + \#(x_j \rightarrow x_i)}
\end{align}
Finally, the probability of generating an ordering $\vec{x} = x_1, x_2, \ldots, x_m$ is computed assuming a first-order Markov chain such that $p(\vec{x}) = \prod_{i=2}^{m} p(x_i|x_{i-1})$.
The \textbf{Probabilistic} ordering strategy enables us to incorporate external knowledge from ConceptNet for the purpose of determining the ordering of input concepts.
Moreover, unlike the \textbf{Example} ordering, \textbf{Probabilistic} ordering can be used both at training and test times.

\section{Experiments}
\label{sec:exp}
\begin{table*}[ht]
\centering
\resizebox{0.95\textwidth}{!}{%
\begin{tabular}{@{}l|l|rrrrrrrr@{}}
\toprule
 & \textbf{Model} & \textbf{BLEU3} & \textbf{BLEU4} & \textbf{ROUGE-2} & \textbf{ROUGE-L} & \textbf{METEOR} & \textbf{CIDEr} & \textbf{SPICE} & \textbf{Coverage} \\ 
\midrule
\multicolumn{1}{c|}{\multirow{4}{*}{\parbox{3cm}{\centering\textbf{Original Ordering~\cite{CommonGen}} \\}}} 
& BERT-Gen & 30.4 & 21.1 & 18.1 & 40.5 & 27.3 & 12.5 & 27.3 & 86.1 \\
\multicolumn{1}{c|}{} & T5-base~ & 26.0 & 16.4 & 14.6 & 34.6 & 23.0 & 9.2 & 22.0 & 76.7 \\
\multicolumn{1}{c|}{} & BART-large~ & 36.3 & 26.3 & \textit{22.2} & 42.0 & \textit{30.9} & 13.9 & 30.6 & \textit{97.4} \\
\multicolumn{1}{c|}{} & T5-large~ & \textit{39.0} & \textit{28.6} & 22.0 & \textit{43.0} & 30.1 & \textit{15.0} & \textit{31.6} & 95.3 \\ 
 \midrule
\multirow{5}{*}{\parbox{3cm}{\centering\textbf{Random} \\ \textbf{ordering}}} 
 & BERT-Gen & 30.5 & 21.1 & 18.0 & 40.1 & 27.4 & 12.6 & 27.4 & 86.5 \\
 & BART-base & 36.0 & 26.6 & 22.0 & 43.3 & 28.8 & 13.9 & 28.5 & 92.0 \\
 & T5-base & 37.7 & 26.9 & 20.9 & 43.2 & 30.2 & 15.1 & 31.2 & 94.2 \\
 & BART-large & \textit{42.4} & \textit{31.8} & \textit{23.8} & 44.7 & \textit{32.8} & \textit{16.6} & \textit{32.3} & \textit{98.8} \\
 & T5-large & 40.6 & 30.0 & 23.1 & \textit{44.9} & 31.5 & 15.9 & 32.1 & 97.6 \\ 
 \midrule
\multirow{5}{*}{\parbox{3cm}{\centering\textbf{Probabilistic} \\ \textbf{ordering}}} 
 & BERT-Gen & 29.1 & 19.9 & 16.9 & 38.6 & 26.7 & 12.2  & 26.7 & 85.7 \\
 & BART-base & 33.5 & 23.8 & 20.4 & 41.2 & 28.0 & 13.4 & 27.8 & 92.6 \\
 & T5-base & 37.6 & 26.8 & 20.5 & 42.4 & 30.6 & 15.1 & 31.1 & 96.4 \\
 & BART-large & 38.5 & 28.0 & 21.8 & 42.4 & \textit{31.7} & 15.4 & 32.2 & \textit{98.8} \\
 & T5-large & \textit{40.8} & \textit{29.8} & \textit{22.2} & \textit{43.9} & 31.6 & \textit{16.2} & \textit{32.4} & 95.7 \\ 
 \midrule
\multirow{5}{*}{\parbox{3cm}{\centering\textbf{Example}\\ \textbf{ordering}}} 
 & BERT-Gen & 33.0 & 23.6 & 19.3 & 41.8 & 28.9 & 13.7 & 28.8 & 93.3 \\
 & BART-base & 36.0 & 26.8 & 22.7 & 44.4 & 29.3 & 14.7 & 29.1 & 97.5 \\
 & T5-base & 40.0 & 30.0 & 22.7 & 45.0 & 31.3 & 16.0 & 32.3 & 97.6 \\
 & BART-large & \textit{\textbf{44.3}} & \textit{\textbf{33.8}} & \textit{\textbf{24.5}} & \textit{\textbf{45.6}} & \textit{\textbf{32.8}} & \textit{\textbf{17.2}} & 33.1 & \textit{\textbf{99.1}} \\
 & T5-large & 43.4 & 32.7 & 23.8 & 45.6 & 32.3 & 16.9 & \textit{\textbf{33.5}} & 98.0 \\ 
 \midrule \midrule
 \multicolumn{1}{l|}{\multirow{5}{*}{\parbox{3cm}{\centering\textbf{Other} \\ \textbf{Prior Methods}}}} 
 & KG-BART~ & 42.1 & 30.9 & 23.4 & 44.5 & 32.9 & 17.5 & 32.7 & 98.7 \\
 & EKI-BART$D_{out}$ & 42.9 & 26.3 & 24.4 & 45.4 & 32.0 & 16.8 & 32.5 & - \\
 & CALM  & - & 29.5 & - & - & 31.9 & 15.6 & 33.2 & - \\
 & NeuroLogic & 41.3 & 36 & - & 44.7 & 31.0 & 15.9 & 31.1 & - \\
 & [MASK] & 43.3 & 32.5 & 24.2 & 44.9 & 32.5 & 17.1 & 32.8 & - \\ 
 \bottomrule
\end{tabular}
}
\caption{Evaluating the quality of the generated sentences on CommonGen test instances.
The input concept sets are ordered using the three strategies described in \autoref{sec:order-methods} during fine-tuning different LMs. Best performing LM fine-tuned with each ordering strategy is shown in italics, while the overall best is shown in bold.}
\label{tab:mainTable}
\end{table*}

\subsection{Experiments setting}
\label{sec:settings}
\paragraph{Dataset:} Our experiments are conducted on the CommonGen dataset~\cite{CommonGen}, which contains 3.5K distinct concept sets (32651/993/1497) with 67389/4018/6042 human written sentences in the training/development/test splits. 
Each instance contains ca. 3-5  input concepts with multiple human-written reference sentences. 

\paragraph{Evaluation Metrics:} We conduct two types of evaluations. First, to evaluate the quality of the generated sentences, we employ the following generation evaluation metrics: BLEU~\cite{bleu}, , ROUGE~\cite{rouge}, METEOR~\cite{meteor}, CIDEr~\cite{cider} and SPICE~\cite{spice}. We report the Coverage of concepts~\cite{CommonGen}, which is defined as the average percentage of input concepts that are present in the generated sentences. Second, we evaluate the ordering of the concepts produced by different LMs (in the order that they appear in the generated sentences) by comparing that to the ordering in human-written test example sentences using the Kendall rank correlation coefficient ($\tau \in [-1,1]$). 
A higher $\tau$ indicates a closer correlation between a concept ordering and their order observed in reference sentences. 
If there are multiple human-written sentences with different orderings for the same input concept set, we take the highest $\tau$ over all of the orderings.

\subsection{Results}
\label{sec:result}
\paragraph{Main Results:} 
We use each concept ordering method to fine-tune five LMs: BERT-Gen~\cite{unilmv2:2020}, T5-base/large~\cite{T5:2019}, and BART-base/large~\cite{BART:2019}.
Implementation details and hyperparameters can be found in \autoref{sec:hyperparameters}.
To assess the effectiveness of the concept set ordering strategies, we compare their performance against two types of prior methods.
The first uses the \textbf{Original} ordering specified in CommonGen~\citep{CommonGen} to fine-tune LMs. 
The second set includes models that incorporate external knowledge, such as KG-BART~\cite{kgbart:2021} and EKI-BART$D_{out}$~\cite{Ekibart:2020}, as well as models that improve performance through pre-training tasks, such as CALM~\cite{CALM:2021}, NeuroLogic~\cite{Neurologic:2021}, and the [MASK]~\cite{Yang:2023}. 

As shown in \autoref{tab:mainTable}, both \textbf{Probabilistic}  and \textbf{Example} ordering outperform the \textbf{Original} Ordering presented in CommonGen~\cite{CommonGen}. 
This shows that refining the ordering strategy for input concepts could enhance an LM's performance in commonsense generation. 
Moreover, \textbf{Example} ordering reports the best performance for all five LMs, even outperforming prior methods that use pre-training tasks or external resources. 

\begin{table}[ht]
\centering
\resizebox{\columnwidth}{!}{%
\begin{tabular}{@{}l|rrrrrr@{}}
\toprule
Ordering & $\tau$ & \textbf{BLEU4} & \textbf{ROUGE-L} & \textbf{METEOR} & \textbf{CIDEr} & \textbf{SPICE} \\ 
 \midrule
Original & 0.328 & 19.0 & 36.4 & 29.0 & 12.9 & 27.7 \\
Random & 0.327 & 18.9 & 36.2 & 29.0 & 12.7 & 27.7 \\
Probabilistic & 0.402 & 20.9 & 37.9 & 29.6 & 13.4 & 29.1 \\ 
\midrule
BERT-Gen & 0.595 & 29.0 & 43.1 & 31.4 & 16.0 & 31.3 \\
BART-base & 0.648 & 31.5 & 44.2 & 31.9 & 16.6 & 32.1 \\
T5-base & 0.627 & 30.3 & 43.9 & 31.9 & 16.4 & 32.0 \\
BART-large & \textbf{0.697} & \textbf{33.4} & \textbf{45.5} & \textbf{32.7} & \textbf{17.3} & \textbf{33.0} \\
T5-large & 0.696 & 32.9 & 45.2 & 32.6 & 17.0 & 32.7 \\
\midrule
Reference & 1 & 35.3 & 57.4 & 33.3 & 18.1 & 33.5\\
\bottomrule
\end{tabular}
}
\caption{Kendall's $\tau$ between the reference ordering and the concept orderings produced using different methods: Lines 2-4 show Original, Random and Probabilistic orderings (\autoref{sec:order-methods}), Lines 5-9 show the orderings of concepts as they appear in the sentences generated using LMs fine-tuned on CommonGen, and Line 10 is the Reference ordering (i.e., ordering of the concepts in human-written CommonGen test sentences).}
\label{tab:diffConceptOrdering}
\end{table}

\begin{table*}[t]
\centering
\resizebox{\textwidth}{!}{%
\begin{tabular}{@{}ll@{}}
\toprule
concept words & throw daughter stream rock daddy \\ \midrule
Reference & Daddy and daughter throwing rocks into stream \\
Turbo(zero shot) & As we hiked beside the stream, my daughter asked if she could throw a rock in the water, but I reminded her that it was too dangerous and daddy \\
Babbage(Original ordering) & little girl throwing a tantrum in the park with her father and sister by the stream and crying because she did not get a toy in the stream. \\
Curie(Original ordering) & mother throwing her daughter in the stream near the rocks with daddy \\
Babbage(Example ordering) & A father and daughter are throwing rocks into a stream. \\
Curie(Example ordering) & daddy and daughter throwing rocks into a stream \\ \midrule
concept words & dog throw frisbee catch \\ \midrule
Reference & A man throws away his dog's favorite frisbee expecting him to catch it in the air. \\
Turbo(zero shot) & The dog loves to play fetch and is always ready to catch the frisbee when it's thrown. \\
Babbage(Original ordering) & A dog is throwing a frisbee. \\
Curie(Original ordering) & A dog is throwing a frisbee and catching it. \\
Babbage(Example ordering) & A dog catching a frisbee and throwing it. \\
Curie(Example ordering) & A dog catching a frisbee thrown by a man. \\ \midrule
concept words & hang squeeze shut head eye \\ \midrule
Reference & Hanging her head, someone squeezes her eyes shut. \\
Turbo(zero shot) & I tried to hang the picture but couldn't do it with just one hand, so I had to squeeze my eyes shut and use both hands to get it done \\
Babbage(Original ordering) & Someone grabs someone by the eyes and squeezes them shut, then drags him away. \\
Curie(Original ordering) & A woman is hanging upside down and squeezing her legs shut, then she opens her eyes. \\
Babbage(Example ordering) & Someone squeezes his eyes shut and hangs his head. \\
Curie(Example ordering) & Someone squeezes his eyes shut and hangs his head. \\ \midrule
concept words & mirror gear picture hold take \\ \midrule
Reference & The man holds the gear and uses the picture taken by the mirror. \\
Turbo(zero shot) & She carefully held the picture up to the mirror to take a closer look at the intricate gear design. \\
Babbage(Original ordering) & A picture of a man in a space suit, holding a mirror, and taking off his gear. \\
Curie(Original ordering) & The mirror has gears in it and a picture of a train engine held up. \\
Babbage(Example ordering) & A man is holding on to a railing to take a picture of his gear and mirror. \\
Curie(Example ordering) & A man is holding a camera and taking a picture of himself in the mirror. \\ \bottomrule
\end{tabular}
}
\caption{Generated examples by different LLMs}
\label{tab:LLMSample}
\end{table*}

\paragraph{Effect of concept ordering:} 
The ordering of concepts as they appear in the human-written CommonGen test sentences (here onwards referred to as the \textbf{Reference} ordering) can be considered as a gold standard for concept ordering.
To evaluate the level of re-ordering of input concepts conducted by an LM, we compare the ordering of concepts in a sentence generated using an LM against their \textbf{Reference} ordering using the Kendall's $\tau$.
Specifically, we first fine-tune each LM using the concept sets as ordered in the CommonGen train instances (i.e., \textbf{Original} ordering).
Next, we present each fine-tuned LM the concept sets in CommonGen test instances in their \textbf{Original} ordering and generate a sentence containing all of the input concepts.
Finally, we compare the ordering of the concepts in the generated sentence against their \textbf{Reference} ordering using $\tau$. The reference sentences in the CommonGen dataset are written by human annotators given the shuffled concepts. Therefore, we think it could be  considered as a “golden” ordering of input concepts. 
To further evaluate the correlation between the ordering of input concepts and the quality of the sentences generated by an LM, we use \texttt{BART-large} fine-tuned with \textbf{Example} ordering as a sentence generator, where we provide it a set of concepts ordered using different methods.


As seen from \autoref{tab:diffConceptOrdering}, among the three concept ordering strategies described in \autoref{sec:order-methods}, the \textbf{Probabilistic} ordering reports the highest $\tau$ value.
In particular, \textbf{Original} has a low $\tau$ value, comparable to \textbf{Random} ordering, indicating that human annotators had to significantly re-order the concept sets shown to them when writing natural sentences.
Interestingly, all five LMs fine-tuned with the \textbf{Example} ordering strategy outperform those three strategies, suggesting that the sentences carry useful contextual clues for determining the ordering of concepts that are exploited by the LMs. We find that the extracted concept ordering generated by BART-large model outperforms that generated by other pre-trained models. It is consistent with the conclusions of \citet{Ou:2022} that a large amount of dependency structure knowledge exists in BART. 
Moreover, we see that $\tau$ values closely aligns with other automatic evaluation metrics, indicating that concept ordering with a higher $\tau$ results in high quality sentence generations. 
Overall, these results suggest that if the concept orderings more closely align with the orderings found in the dataset reference sentences, the Example Ordering trained model could generate superior sentences, leveraging the inherent commonsense knowledge embedded within the pre-trained models.

\subsection{Large Language Model Scenario}
\label{sec:LLM}
Large Language Models (LLMs) have shown impressive performance in complex reasoning compared to the smaller size models~\cite{GPT3:2020, lamda:2022}. 
To investigate their effectiveness for concept ordering, we use the GPT-3 and GPT-3.5~\cite{GPT3:2020} as LLMs and use a prompt-based instruction (see \autoref{sec:LLMdetails} for the details of the prompts) to induce an ordering among a given set of concepts. 
Due to the space limitations we show the results in the Appendix \autoref{tab:LLMEvaluation}.
We find that the Example Ordering outperforms the unordered inputs, indicating that concept ordering is also important to the LLMs. 
Moreover, from the generated sentences (shown in \autoref{tab:LLMSample}) we find that the Example ordering improves their quality as well as the concept coverage. 
Interestingly however, the best results obtained using LLMs with prompts are worse than that compared to the fine-tuned \texttt{BART-large}, which indicates importance of fine-tuning on CommonGen. 

\section{Conclusion}
\label{sec:conclusion}
We examined the impact of concept ordering on the quality of generated sentences in GCR using multiple LMs.
We find that ordering the input concepts can improve performance, and all fine-tuned LMs generate better quality sentences when the input concepts were presented in an order consistent with that found in human-written sentences. 

\section{Limitations}
\label{sec:limit}
In this work, we limited our investigation to the generation of English sentences and to a finite set of pre-trained language models (PLMs).  
This limitation was largely due to the nature of the CommonGen dataset, the only publicly available dataset we found for concept-to-sentence tasks, which is primarily English-centric. 
Therefore, our evaluation of the generation quality was limited to English, which is a morphologically limited language.
Different languages have different grammars and sentence structures, but the automatic evaluation metrics such as BLEU~\cite{bleu}, ROUGE~\cite{rouge} and CIDEr~\cite{cider} could also be used for other languages. 
Therefore, we consider it to be an important next step to evaluate the concept ordering strategies described in this work on languages other than English.

Furthermore, we focus only on the experiments around concept ordering in the LLM scenario. 
However, as reported in much prior work, the commonsense knowledge inside the LLMs could be efficiently exploited by intermediate reasoning steps and designed prompts~\cite{wei:2022,zhang:2023}.  We utilised identical prompts for each backbone LLM and did not investigate the potential influence of prompt variations. we observed that some sentences generated by LLMs contradict commonsense. Therefore, generative commonsense reasoning tasks remain a challenge for LLMs. Given the scope of the current paper, research on prompt design and encoding concept ordering into intermediate steps will be pursued in future work.

Given a set of concepts, it is possible to write multiple natural sentences that arrange the input concepts in different orders.
However, in the CommonGen dataset only a small number of human-written sentences are available for a given concept set, which does not cover all possible orderings.
Therefore, the Kendall's $\tau$ values reported in this paper must be taken as a lower bound on the agreement with a human determining the ordering of concepts in a sentence.
Moreover, we it is important to consider other types of rank evaluation metrics in addition to Kendall's $\tau$ for future evaluations.

\section{Ethical Considerations}
\label{ethical}

While we conducted our research primarily on the CommonGen dataset, which to the best of our knowledge does not present any explicit ethical issues, it is essential to acknowledge the potential for social biases in the LMs~\cite{blodgett2021stereotyping}.
One of the pre-trained language models we used in our experiments, BART are significantly prone to prediction errors related to gender bias~\cite{sharma:2021} and we are not evaluating for the biases in the generated sentences here which should be done before LMs are deployed in the downstream NLP applications. 
Given that we are fine-tuning LMs on the CommonGen dataset, some social biases could get amplified during this fine-tuning process.
The predicted sentences are possibly influenced by such biases. 
It is still an open question for how to effectively mitigate these biases, particularly in the context of generative commonsense reasoning tasks.

\bibliography{GCR-order.bib}
\bibliographystyle{acl_natbib}

\appendix
\section*{Supplementary Appendix}
\begin{table*}[ht]
\centering
\begin{tabular}{l|rrrrrrr}
\toprule
 & \textbf{Parameters} & \textbf{BLEU4} &  \textbf{ROUGE-L} & \textbf{METEOR} & \textbf{CIDEr} & \textbf{SPICE}  \\ 
\midrule
Turbo(zero shot) & 135B & 13.9 &  33.5 & 27.2 & 8.7 & 24.0  \\ 
\midrule
Babbage (Original) & 6.7B & 19.7 &  37.7 & 27.1 & 12.0 & 26.7  \\
Curie(Original) & 13B  & 13.5 &  32.6 & 26 & 10.0 & 24.6  \\ 
\midrule
Babbage (Example) & 6.7B  & 23.6 &  41.0 & 29.2 & 13.6 & 29.2  \\
Curie (Example) & 13B  & 25.5 &  41.9 & 30.0 & 14.4 & 30.1  \\
\bottomrule
\end{tabular}
\caption{Evaluation of the effect of different concept ordering strategies on the LLMs. The results show that a good concept ordering strategy could also help the quality of large model generation.}
\label{tab:LLMEvaluation}
\end{table*}
\section{Different Input Formats for Example Ordering}
\label{sec:format}
We conducted an evaluation to examine the impact of input formats on the quality of generation. 
We designed three types of input formats, given a concept set: 
1) the ordered concepts are concatenated together, 
2) the ordered concepts are concatenated together and separated (delimitted) by commas, 
and 3) the unordered concepts are concatenated with the ordered concepts and denoted by a special ``\textit{[ORDERING]}'' tokens. 
An example using the concept set {\emph{ski}, \emph{mountain}, \emph{skier}}, is provided in \autoref{tab:input_examples}. 
As shown in \autoref{tab:diff_inputs}, the performance of the generated sentences does not change significantly with the different input formatting methods.
However, the best performance among the three input formatting methods is achieved when the concepts are concatenated with the ordered concepts around tokens. 
This may be attributed to the fact that, with the tokens and ordered concepts, a model might better comprehend the ordering task and use the syntactic knowledge embedded in the pre-trained models~\cite{Ou:2022}.

\begin{table}[h]
\centering
\begin{tabular}{@{}l@{\hspace{0.5cm}}l@{}}
\toprule
\textbf{Format} & \textbf{Example} \\
\midrule
w/o Comma & skier ski mountain \\
Comma & skier, ski, mountain \\
Token & ski mountain skier \texttt{[ORDERING]} \\
& skier ski mountain \texttt{[ORDERING]} \\
\bottomrule
\end{tabular}
\caption{Three different input formats given a concept set \{\emph{ski}, \emph{mountain}, \emph{skier}\}}
\label{tab:input_examples}
\end{table}

\begin{table}[h]
\centering 
\resizebox{\columnwidth}{!}{%
\begin{tabular}{@{}l|rrrrr@{}}
\toprule
\textbf{Concept set} & \textbf{BLEU4} & \textbf{ROUGE-L} & \textbf{METEOR} & \textbf{CIDEr} & \textbf{SPICE}  \\ 
\midrule
w/o Comma &  33.4 & 45.5 & 32.6 & 17.1 & 32.8  \\
Comma &  32.5 & 45 & 32.4 & 16.8 & 32.6  \\
Token &  33.8 &  45.6 & 32.8 & 17.2 & 33.1  \\ 
\bottomrule
\end{tabular}
}
\caption{The generated evaluation of different inputs formats.}
\label{tab:diff_inputs}
\end{table}

\section{Concept Ordering on LLMs}
\label{sec:LLMdetails}
We also conducted concept ordering experiments using the CommonGen dataset on LLMs containing more than one billion parameters. 
Specifically, we selected GPT3 models (Babbage and Curie) and GPT3.5 Turbo from OpenAI as our backbone models. 
Since the Turbo model does not support fine-tuning, it was evaluated solely in a zero-shot setting. 
We evaluated the quality of the generated outputs using both the Original ordering and Example ordering strategies on the Babbage and Curie models. 

As shown in \autoref{tab:LLMEvaluation}, the Example Ordering strategy also enhanced the performance of LLMs on the CommonGen dataset task. 
However, we observed that the three LLMs underperform compared against the smaller fine-tuned LMs. 
We notice that this problem also exists in other NLP tasks~\cite{GPT3:2020,gutierrez:2022}. 
Therefore, we examined some sentences generated by these LLMs as presented in \autoref{tab:LLMSample}.
Under the zero-shot setting, Turbo was capable of generating more complex sentences than those provided in the references, and these sentences were consistent with commonsense. When we try to use Example Ordering Strategy for the Turbo model with the prompt \textit{Given a concept list: [concepts], please generate a sentence that aligns the ordering of the concepts:}, we find that the model could not always follow the concept ordering given even under the few-shot setting.

For the Babbage and Curie models, the Original Ordering strategy often resulted in generated sentences that did not encompass all concepts from the given input set. 
The Example Ordering strategy ameliorated this issue; however, even when the grammar was correct, some sentences were inconsistent with commonsense (e.g., \textit{A dog catching a Frisbee and throwing it}). 
We surmise that these issues are the primary reasons why the LLMs did not perform well.
As is suggested by \href{https://platform.openai.com/docs/guides/fine-tuning/preparing-your-dataset}{OpenAI website}, the performance of the model is based on the description given and external content would improve the performance. 
However, the input of CommonGen only has a single concept set. Potential improvements for the performance of LLMs on such tasks could involve providing more content around the concept set such as description about the relations among the concepts.

\section{Training Details}
\label{sec:hyperparameters}
We utilised the BERT, BART, and T5 models as implemented in the Transformers library~\cite{transformers:2020}. Detailed hyper-parameters are provided in the accompanying bash scripts (submitted as supplementary materials).
The primary hyper-parameters were initialised in alignment with the standards set out in the CommonGen paper~\cite{CommonGen}. We fine-tuned each model using the ADAM optimiser~\cite{Adam}. 
To prevent over-fitting, we adopted an early stopping strategy based on the development set's loss. 
We train the model with one NVIDIA A6000 GPU and one V100 GPU. 
Each model could be fine-tuned in less than 6 hours.

The training of the LLMs utilised OpenAI's API.\footnote{\url{https://platform.openai.com/docs/guides}}
Training a Babbage model incurred a cost of \$3 for the CommonGen dataset, while training a Curie model incurred a cost of \$14. During model training, each instance would start with the prompt ``\textit{Generate a sentence containing all the concepts in the concept set:}'' for each concept set and we added a separator `` -\textgreater''  at the end of the prompt. We also use an ending token ``/n'' at the end of the target sentence.
\begin{table}[h]
\centering
\resizebox{\columnwidth}{!}{%
\begin{tabular}{@{}l@{\hspace{0.6cm}}l@{\hspace{0.6cm}}l@{\hspace{0.6cm}}l@{}}
\toprule
\textbf{Models} & \textbf{Random} & \textbf{Probabilistic} & \textbf{Example} \\
\midrule
BERT-Gen & 32*2, 5e-5 & 32*2, 5e-5 & 32*2, 5e-5 \\
BART-base & 64*4, 3e-5 & 64*3, 5e-5 & 64*6, 3e-5 \\
T5-base & 64*4, 5e-5 & 64*3, 5e-5 & 64*4, 5e-5 \\
BART-large & 96*4, 3e-5 & 64*4, 2e-5 & 64*4, 3e-5 \\
T5-large & 32*4, 2e-5 & 64*4, 2e-5 & 64*4, 3e-5 \\
GPT3-babbage & - & - & 128, - \\
GPT3-curie & - & - & 128, - \\
\bottomrule
\end{tabular}%
}
\caption{Key parameters in experiments. In each cell, the left is the batch size and the right is the learning rate.}
\label{tab:keyparameters}
\end{table}


\end{document}